OpenStudy (anonymous):

Consider the function y=x^2+5x-2 . What would happen to the graph if (x + 5) was substituted in place of the x? A. The graph would shift 5 units to the right. B. The graph would shift down 5 units. C. The graph would shift up 5 units. D. The graph would shift 5 units to the left. Consider the function y=x^2+5x-2 . What would happen to the graph if (x + 5) was substituted in place of the x? A. The graph would shift 5 units to the right. B. The graph would shift down 5 units. C. The graph would shift up 5 units. D. The graph would shift 5 units to the left. @Mathematics

OpenStudy (anonymous):

A

OpenStudy (anonymous):

option D

OpenStudy (anonymous):

oopsie its is D

OpenStudy (shaik0124):

EXPLAIN HOW U GOT D ANSWER

OpenStudy (anonymous):

first find roots

OpenStudy (anonymous):

and then find the roots after (x+5) is substitutes

OpenStudy (anonymous):

then it would be obvious that its been moved 5 to the left

OpenStudy (shaik0124):

WHAT EQN DID U GOT

OpenStudy (anonymous):

15x^4-20x^3+10x^2-35x

OpenStudy (anonymous):

thats with x+5 substituted

OpenStudy (anonymous):

x+5=x-(-5) so, the graph will be valid if origin is taken (-5,0) instead of (0,0) so, the whole graph will shift in 5 units to left

OpenStudy (anonymous):

it actually x^2+15x+48

OpenStudy (anonymous):

after x+5 is substituted

OpenStudy (shaik0124):

I GOT THE EQN X^2+15X-47

OpenStudy (anonymous):

no its +48

OpenStudy (anonymous):

not -47

OpenStudy (shaik0124):

YA =48

OpenStudy (anonymous):

why u guys r fighting on it?? :)

OpenStudy (anonymous):

Which of the following describes the non-rigid transformation in the function shown below? y+5=-2(x-1)^2 A. The graph is reflected across the x-axis. B. The graph is shifted 1 unit right. C. The graph is shifted 5 units down. D. The graph is stretched vertically by a factor of 2.

OpenStudy (anonymous):

im just saying the right answer not fighting

OpenStudy (anonymous):

ok, ok :D

OpenStudy (anonymous):

but finding the roots- method is not so good, i guess

OpenStudy (anonymous):

is there a better way?

OpenStudy (anonymous):

hmm, think it this way in the first case, the result u will get substituting x=0, in the 2nd case u will get that exact value for x=-5

OpenStudy (anonymous):

oh ok